# Real-time Tracking in a Status Update System with an Imperfect Feedback Channel

### v1

@@ V Conclusions
We considered an energy harvesting status update system monitoring a finite-state Markov source where both forward and feedback channels are error-prone. We addressed the real-time tracking problem by developing policies that minimize the long-term time average of a distortion function subject to an energy causality constraint. We modeled the stochastic problem as a POMDP and then cast it as a belief-MDP problem. By truncating the belief-state space, we proposed a transmission policy using RVIA. Furthermore, we proposed a sub-optimal low-complexity transmission policy by transforming the belief-MDP problem into a sequence of per-slot problems. We numerically examined the performance and the structure of the proposed policies. The results show that the RVIA-based policy has a switching-type structure. Moreover, we observed that when the channel condition is good and the source state changes more frequently, 1) the proposed RVIA-based policy decides to transmit status updates less frequently (for $p=q$ it decides to stay idle) and 2) the low-complexity policy performs close to the and RVIA-based policy.
@@ Appendix A Proof of Proposition 1
To prove that remaining in the most recently received successful status update is the optimal ML estimator for the sink when $p>q$, we need to show that after $K$ transitions, the probability of self-transition is larger or equal to the probability of transition to the other states. We use the induction approach. We first show that it is valid for one transition. Then, considering that it holds for ${K-1}$ transitions, we prove it holds for ${K}$ transitions. The one transition $P^{2}$ is given as
$\begin{array}[]{ll}P^{2}=P\times P=\begin{bmatrix}R_{1,1}&R_{1,2}&\dots&&&R_{1,2}\\ R_{1,2}&R_{1,1}&&&&\vdots\\ \vdots&&\ddots&&&R_{1,2}\\ R_{1,2}&\dots&R_{1,2}&&&R_{1,1}\end{bmatrix},\end{array}$ (28)
where ${R_{1,1}=p^{2}+(N-1)q^{2}}$ and ${R_{1,2}=2pq+(N-2)q^{2}}$. Now, we calculate $R_{1,1}-R_{1,2}$ to determine which one is larger.
@@
Now, we assume the probability of self-transition is larger and equal to the transition to other states after $K-1$ transition and show it holds for $K$ transition $P^{K+1}$ given as
@@
where ${R_{K,1}=pR_{K-1,1}+(N-1)qR_{K-1,2}}$ and ${R_{K,2}=pR_{K,2}+qR_{K-1,1}+(N-2)qR_{K-1,2}}$. Now, Now, we calculate $R_{K,1}-R_{K,2}$ to determine which one is larger.
@@ Appendix B Derivation of the Conditional Probabilities in (10)
$\mathrm{Pr}(E_{1}\mid\overline{\text{ACK}})=\frac{\mathrm{Pr}(E_{1}\cap\overline{\text{ACK}})}{\mathrm{Pr}(\overline{\text{ACK}})}$
=$\hskip 34.1433pt\overset{(a)}{=}\frac{\mathrm{Pr}(E_{1})}{\mathrm{Pr}(\overline{\text{ACK}})}$
=$\hskip 34.1433pt\overset{(b)}{=}\frac{p_{s}(1-p_{f})}{\mathrm{Pr}(\overline{\text{ACK}})}$
=$\hskip 34.1433pt\overset{(c)}{=}\frac{p_{s}(1-p_{f})}{1-p_{s}p_{f}},$ (32)
@@
$\mathrm{Pr}({E_{2}}\mid\overline{\text{ACK}})$ $=\frac{\mathrm{Pr}(E_{2}\cap\overline{\text{ACK}})}{\mathrm{Pr}(\overline{\text{ACK}})}$
=$\overset{(a)}{=}\frac{\mathrm{Pr}(E_{2})}{\mathrm{Pr}(\overline{\text{ACK}})}$
@@ Appendix C Proof of Theorem 1
To show the belief-MDP is communicating, it is sufficient to find a randomized policy that induces a recurrent Markov chain, i.e., a policy under which any belief-state ${l^{\prime}=(X^{\prime},b^{\prime},\rho^{\prime},X_{-1}^{\prime},d_{-1}^{\prime})\in\mathcal{L}}$ can be reached from any other state ${l=(X,b,\rho,X_{-1},d_{-1})\in\mathcal{L}}$ with a positive probability [34, Prop. 8.3.1]. We define the following policy: the transmitter sends a status update, i.e., $a(t)=1$, with probability $0.5$ whenever the battery is not empty, i.e.,
@@
We consider two cases $b^{\prime}\geq b$ and $b^{\prime}<b$. For the case where $b^{\prime}\geq b$, let us consider the following event: for $b^{\prime}-b$ consecutive slots the transmitter stays idle (i.e., $a=0$) and it receives energy during each slot (i.e., $e=1$), at the next slot, the source state moves to the state regarding the reset belief associated with $\rho^{\prime}$, the transmitter sends an update (i.e., $a=1$) and receives ACK, and depending on the values of $d_{-1}^{\prime}$ and $\rho^{\prime}$ we consider the following three cases: i) if $d_{-1}^{\prime}=1$ and $\rho^{\prime}$ is a reset belief, at the next slot, the source’s state moves to $X^{\prime}$, and the transmitter receives one unit of energy, ii) if $d_{-1}^{\prime}=0$ and $\rho^{\prime}$ is a reset belief, at the next two slots, first, the source state moves to $X_{-1}^{\prime}$ and then moves to $X^{\prime}$, and the transmitter receives one unit of energy, and iii) if $d_{-1}^{\prime}=0$ and $\rho^{\prime}$ is not a reset belief, until moving to $\rho^{\prime}$, at each slot, the transmitter receives one unit of energy, the transmitter sends an update without receiving ACK, and the source state changes in a way that the belief becomes $\rho^{\prime}$ (see equation (10)), then, at the next two slots, first, the source state moves to $X_{-1}^{\prime}$ and then moves to $X^{\prime}$, and the transmitter receives one unit of energy. Since the considered event occurs with a positive probability, any belief-state ${l^{\prime}=(X^{\prime},b^{\prime},\rho^{\prime},X_{-1}^{\prime},d_{-1}^{\prime})\in\mathcal{L}}$ can be reached from any other state ${l=(X,b,\rho,X_{-1},d_{-1})\in\mathcal{L}}$ with a positive probability for $b^{\prime}\geq b$. For the case $b^{\prime}<b$, first, the battery is discharged to $b^{\prime}$ by sending updates for $b-b^{\prime}$ slots without receiving energy, then a similar event with a positive probability can be provided.
Appendix D The derivation of (III-D)
@@
$\begin{cases}g\big{(}X(t),X(t)\big{)}&E=\big{(}1,1,1\big{)}\\ \displaystyle\sum_{i=1}^{N}g\big{(}X(t),i\big{)}\rho_{i}(t)&E=(1,0,1)\\ \displaystyle\sum_{j=1,j\neq X(t)}^{N}g\big{(}j,X(t)\big{)}&E=(1,1,0)\\ \displaystyle\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}g(j,i)\rho_{i}(t)&E=(1,0,0)\\ \displaystyle\sum_{i=1}^{N}g\big{(}X(t),i\big{)}\rho_{i}(t)&E=(0,0,1)\\ \displaystyle\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}g(j,i)\rho_{i}(t)&E=(0,0,0).\end{cases}$ (35)
@@
$\mathbb{E}\{C(t+1)\mid l(t)\}=g\big{(}X(t),X(t)\big{)}\mathrm{Pr}\big{(}E=(1,1,1)\mid l(t)\big{)}$
$+\sum_{i=1}^{N}g\big{(}X(t),i\big{)}\rho_{i}(t)\mathrm{Pr}\big{(}E=(1,0,1)\mid l(t)\big{)}$
$+\sum_{j=1,j\neq X(t)}^{N}g\big{(}j,X(t)\big{)}\mathrm{Pr}\big{(}E=(1,1,0)\mid l(t)\big{)}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}g(j,i)\rho_{i}(t)\mathrm{Pr}\big{(}E=(1,0,0)|l(t)\big{)}$
$+\sum_{i=1}^{N}g\big{(}X(t),i\big{)}\rho_{i}(t)\mathrm{Pr}\big{(}E=(0,0,1)\mid l(t)\big{)}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}g(j,i)\rho_{i}(t)\mathrm{Pr}\big{(}E=(0,0,0)\mid l(t)\big{)}.$ (36)
@@
$\mathbb{E}\{C(t+1)\mid l(t)\}=g\big{(}X(t),X(t)\big{)}pp_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}g\big{(}X(t),i\big{)}\rho_{i}(t)p(1-p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{j=1,j\neq X(t)}^{N}g\big{(}j,X(t)\big{)}(1-p)p_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}g(j,i)\rho_{i}(t)(1\!-\!p)(1\!-\!p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}g\big{(}X(t),i\big{)}\rho_{i}(t)(1-p)(1-\mathbb{E}\{a(t)\mid l(t)\})$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}g(j,i)\rho_{i}(t)p(1-\mathbb{E}\{a(t)\mid l(t)\}).$ (41)

### v2

@@ V Conclusions
We considered an energy harvesting status update system monitoring a finite-state Markov source where both forward and feedback channels are error-prone. We addressed the real-time tracking problem by developing policies that minimize the long-term time average of a distortion function subject to an energy constraint imposed by the energy harvesting circuit. We modeled the stochastic problem as a POMDP and then cast it as a belief-MDP problem. By truncating the belief-state space, we proposed a transmission policy using RVIA. Furthermore, we proposed an energy-agnostic LC policy by transforming the belief-MDP problem into a sequence of per-slot problems. Then, we extended the energy-agnostic LC policy to an energy-aware LC policy by adding a regularization term to the objective function of the per-slot problems. We numerically examined the performance and structure of the proposed policies. The results showed that the proposed policies have a switching-type structure. The structures suggest that when the source dynamic is high, the optimal action is to send updates less frequently. In addition, we observed that the energy-aware LC policy performs relatively close to the POMDP-based policy. Also, for scenarios where the energy arrival rate is high and the source dynamic is low, the energy-agnostic LC policy performs close to the POMDP-based policy.
@@ Appendix A Proof of Proposition 1
To show that staying in the most recent successful status update is the optimal ML estimator when $p>q$, we must prove that after $K$ transitions, the probability of self-transition is at least as large as the probability of transitioning to other states. We use the induction approach. We first show that it is valid for one transition. Then, considering that it holds for ${K-1}$ transitions, we prove it holds for ${K}$ transitions.
Let $P$ denote the state transition matrix of the source, given as
Note that $P$ is a doubly stochastic matrix, i.e., a square matrix with non-negative entries between 00 and $1$, where each row and column sums up to $1$. The one transition $P^{2}$ is given as
$\begin{array}[]{ll}P^{2}=P\times P=\begin{bmatrix}R_{1,1}&R_{1,2}&\dots&R_{1,2}\\ R_{1,2}&R_{1,1}&&\vdots\\ \vdots&&\ddots&R_{1,2}\\ R_{1,2}&\dots&R_{1,2}&R_{1,1}\end{bmatrix},\end{array}$ (27)
where ${R_{1,1}=p^{2}+(N-1)q^{2}}$ and ${R_{1,2}=2pq+(N-2)q^{2}}$. Now, we calculate $R_{1,1}-R_{1,2}$
@@
Now, we assume the probability of self-transition is larger and equal to the transition to other states after $K-1$ transition and show it holds for $K$ transitions. The $K$ transitions $P^{K+1}$ is given as
@@
where ${R_{K,1}=pR_{K-1,1}+(N-1)qR_{K-1,2}}$ and ${R_{K,2}=pR_{K,2}+qR_{K-1,1}+(N-2)qR_{K-1,2}}$. Now, we calculate $R_{K,1}-R_{K,2}$ to determine which one is larger.
@@ Appendix B Derivation of the Conditional Probabilities in (8)
$\mathrm{Pr}(E_{1}\mid\overline{\text{ACK}})=\frac{\mathrm{Pr}(E_{1}\cap\overline{\text{ACK}})}{\mathrm{Pr}(\overline{\text{ACK}})}\overset{(a)}{=}\frac{\mathrm{Pr}(E_{1})}{\mathrm{Pr}(\overline{\text{ACK}})}$
=$\hskip 34.1433pt\overset{(b)}{=}\frac{p_{s}(1-p_{f})}{\mathrm{Pr}(\overline{\text{ACK}})}\overset{(c)}{=}\frac{p_{s}(1-p_{f})}{1-p_{s}p_{f}},$ (29)
@@
$\mathrm{Pr}({E_{2}}\mid\overline{\text{ACK}})$ $=\frac{\mathrm{Pr}(E_{2}\cap\overline{\text{ACK}})}{\mathrm{Pr}(\overline{\text{ACK}})}\overset{(a)}{=}\frac{\mathrm{Pr}(E_{2})}{\mathrm{Pr}(\overline{\text{ACK}})}$
@@ Appendix C Proof of Theorem 1
To show the belief-MDP is communicating, it is sufficient to find a randomized policy that induces a recurrent Markov chain, i.e., a policy under which any belief-state ${l^{\prime}=(X^{\prime},b^{\prime},\rho^{\prime},X_{-1}^{\prime},f_{-1}^{\prime})\in\mathcal{L}}$ can be reached from any other state ${l=(X,b,\rho,X_{-1},f_{-1})\in\mathcal{L}}$ with a positive probability [38, Prop. 8.3.1]. We define the following policy: the transmitter sends a status update, i.e., $a(t)=1$, with probability $0.5$ whenever the battery is not empty, i.e.,
@@
We consider two cases $b^{\prime}\geq b$ and $b^{\prime}<b$. For the case where $b^{\prime}\geq b$, let us consider the following event: for $b^{\prime}-b$ consecutive slots the transmitter stays idle (i.e., $a=0$) and it receives energy during each slot (i.e., $e=1$), at the next slot, the source state moves to the state regarding the reset belief associated with $\rho^{\prime}$, the transmitter sends an update (i.e., $a=1$) and receives ACK, and depending on the values of $f_{-1}^{\prime}$ and $\rho^{\prime}$ we consider the following three cases: i) if $f_{-1}^{\prime}=1$ and $\rho^{\prime}$ is a reset belief, at the next slot, the source’s state moves to $X^{\prime}$, and the transmitter receives one unit of energy, ii) if $f_{-1}^{\prime}=0$ and $\rho^{\prime}$ is a reset belief, at the next two slots, first, the source state moves to $X_{-1}^{\prime}$ and then moves to $X^{\prime}$, and the transmitter receives one unit of energy, and iii) if $f_{-1}^{\prime}=0$ and $\rho^{\prime}$ is not a reset belief, until moving to $\rho^{\prime}$, at each slot, the transmitter receives one unit of energy, the transmitter sends an update without receiving ACK, and the source state changes in a way that the belief becomes $\rho^{\prime}$ (see equation (8)), then, at the next two slots, first, the source state moves to $X_{-1}^{\prime}$ and then moves to $X^{\prime}$, and the transmitter receives one unit of energy. Since the considered event occurs with a positive probability, any belief-state ${l^{\prime}=(X^{\prime},b^{\prime},\rho^{\prime},X_{-1}^{\prime},f_{-1}^{\prime})\in\mathcal{L}}$ can be reached from any other state ${l=(X,b,\rho,X_{-1},f_{-1})\in\mathcal{L}}$ with a positive probability for $b^{\prime}\geq b$. For the case $b^{\prime}<b$, first, the battery is discharged to $b^{\prime}$ by sending updates for $b-b^{\prime}$ slots without receiving energy, then a similar event with a positive probability can be provided.
Appendix D The derivation of (III-D1)
@@
$\begin{cases}d\big{(}X(t),X(t)\big{)}&E=\big{(}1,1,1\big{)}\\ \displaystyle\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)&E=(1,0,1)\\ \displaystyle\sum_{j=1,j\neq X(t)}^{N}d\big{(}j,X(t)\big{)}&E=(1,1,0)\\ \displaystyle\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)&E=(1,0,0)\\ \displaystyle\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)&E=(0,0,1)\\ \displaystyle\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)&E=(0,0,0).\end{cases}$ (31)
@@
$\mathbb{E}\{C(t+1)\mid l(t)\}=d\big{(}X(t),X(t)\big{)}\mathrm{Pr}\big{(}E=(1,1,1)\mid l(t)\big{)}$
$+\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)\mathrm{Pr}\big{(}E=(1,0,1)\mid l(t)\big{)}$
$+\sum_{j=1,j\neq X(t)}^{N}d\big{(}j,X(t)\big{)}\mathrm{Pr}\big{(}E=(1,1,0)\mid l(t)\big{)}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)\mathrm{Pr}\big{(}E=(1,0,0)|l(t)\big{)}$
$+\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)\mathrm{Pr}\big{(}E=(0,0,1)\mid l(t)\big{)}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)\mathrm{Pr}\big{(}E=(0,0,0)\mid l(t)\big{)}.$ (32)
@@
$\mathbb{E}\{C(t+1)\mid l(t)\}=d\big{(}X(t),X(t)\big{)}pp_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)p(1-p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{j=1,j\neq X(t)}^{N}d\big{(}j,X(t)\big{)}qp_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)q(1\!-\!p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)p(1-\mathbb{E}\{a(t)\mid l(t)\})$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)q(1-\mathbb{E}\{a(t)\mid l(t)\}).$